OpenStudy (anonymous):

√5:(x^2) in exponential form

OpenStudy (turingtest):

is this the problem?\[\sqrt{5x^2}\]

OpenStudy (anonymous):

no

OpenStudy (turingtest):

can you try to type it with the equation feature then?

OpenStudy (anonymous):

^5√x^2

OpenStudy (turingtest):

oh you mean\[\sqrt[5]{x^2}\]right?

OpenStudy (anonymous):

yes ;)

OpenStudy (turingtest):

The rule in general is\[\sqrt[b]{x^a}=x^{a/b}\]so in your case a=2 b=5 which gives us\[\sqrt[5]{x^2}=x^{2/5}\]

OpenStudy (turingtest):

To type these in the future write it like this: sqrt[5]{x^2} Then highlight it and click on the "equation" feature (in the lower left). You should get used to using it so we can read your problems correctly.

OpenStudy (anonymous):

49(-1/2) Simplify

OpenStudy (turingtest):

there is not much to be done here because 49 is not divisible by 2, so just multiply across:\[49*{-1\over2}={49\over1}*{-1\over2}=?\]you should be able to do that yourself.

OpenStudy (anonymous):

49(-1/2) Simplify A. -7 B. 7 C, -1/7 D. 1/7

OpenStudy (turingtest):

oh you wanted \[49^{-1/2}=\sqrt{49^{-1}}=\sqrt{1\over49}={1\over \sqrt{49}}=\pm{1\over7}\]so both C and D are correct

OpenStudy (turingtest):

please use ^ symbol for exponents: 49^(-1/2) is how your problem should be written.